# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 3.2 Background: Error detection and correction
A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^6] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can readily compute (cf. Sec. 23.4). while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large. We can compute the coefficients of the wild factor $q(x):=p(x)/\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $q(x)$ the same or another root-finder, and repeat this computation recursively until we approximate all $m$ roots in the disc. [^7] By this recipe we detect output errors at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them. Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them.
Readily bounds the complexity and error probability of this recursive process:
@@
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}.$ (3.3)
One can accelerate low precision numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) of [50, 28, 32].
3.3 Cauchy-based root-counting and e/i tests
3.3.1 Introductory comments
Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
[…]
Having an approximation of the integral of (3.4) within less than 1/2, we can round it to the closest integer and obtain $s_{0}=\#(D(0,1))$. This is more than enough for our e/i tests, for which we only need to know whether $s_{0}=0$ or not.
3.3.2 Cauchy-based soft e/i tests: outline
In Part II we approximate the integral of (3.4) with finite sums made up of the scaled values of NIR$(x)$ at the scaled $q$th roots of unity. Namely, for a fixed $q>1$ let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity:
$\zeta:=\zeta_{q}:=\exp\Big{(}\frac{2\pi{\bf i}}{q}\Big{)},~{}{\bf i}:=\sqrt{-1},$ (3.5)
and devise randomized $\sigma$-soft e/i tests as follows:
(i) Choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$.
(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
[…]
(iv) Conclude that exclusion holds whp if $|s_{0,q}|<1/2$.
(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$, from $\frac{1}{\gamma}$ to $\frac{1}{\gamma^{v}}$ for $\gamma>1$ of our choice, say, for $\gamma=2$.
The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points.
3.4 Soft e/i test with root-lifting: generic algorithm
In Part III we decrease the complexity estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer is 0 or not, which is sufficient for e/i tests. In these e/i tests we elaborate upon the generic one (see Alg. 3.1 below), where we use $h$-lifting of roots:
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi{\bf i}}{h}\Big{)},$ (3.6)
@@ 3.4 Soft e/i test with root-lifting: generic algorithm
The $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [51]):
@@
Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.7) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.3) from any complex point (see [86, 113, 115] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [115], but even the straightforward implementation of iterations (3.7) promises to be highly efficient as well.
We will elaborate upon the following generic algorithm in Sec. 14.
@@
INITIALIZATION: Fix a positive $q\in\mathbb{N}$ and a positive integer $h$.
COMPUTATIONS:
(1) Select $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
(2) By extending (3.2) to $p_{h}(x)$ of (3.7) (we specify this extension in (13.3)) approximate the values
@@
(3) If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
@@ 3.5 Soft e/i test with root-lifting: correctness of inclusion claim
It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D(c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
@@
We use Cor. 3.1 later but next apply Cor. 3.2 to prove correctness of the inclusion claim of Alg. 3.1. The bound of Cor. 3.1 on $r_{d}(c)$ is too large for that, but a similar bound on $r_{d}(c,p_{h})=(r_{d}(c))^{h}$, implies that $r_{d}(c)$ is close to 1 already for $h$ of order $\log(d)$.
@@ 3.6 Soft e/i test with root-lifting: further analysis (outline)
(i) Correctness of the exclusion claim of Alg. 3.1. We prove it separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (3.7) or root-lifting (3.6).
(a) In Sec. 14.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Example 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
@@
(ii) Counting query locations. Eqn. (3.7) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all the $h$th roots of $c$. We can fix $h>1$ being a power of 2 and apply (3.7) or fix any integer $h>1$ and apply (3.6). In both cases Alg. 3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion. We always choose $h=O(\log(d))$ and then can ignore a factor of $h$ in our $\tilde{O}$ estimates.
In our deterministic soft e/i test based on Alg. 3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points, combine these estimates with Cor. 2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
Under the Random Root Model of Sec. 1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$. We strengthen this isolation by applying $O(\log(\log(d)))$ root-squaring steps (3.7) or $O(\log(d))$ steps of root-lifting of (3.6) and then whp decrease the overall bit operation count from $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$.
(iii) Estimation of computational precision and bit-operation complexity. Louis and Vempala in [76] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations. We combine a small part of their advanced analysis with Las Vegas randomization and some customary techniques of error analysis to extend this estimate to our root-finders.
For the solution of Problems 0, 1, and 3, where the coefficients are given, we apply fast algorithm of [81] for multipoint polynomial evaluation to decrease the estimated number of queries and the bit operation cost bound.
We devise two groups of efficient e/i tests in Parts II and III of this paper, and in both cases arrive at the estimates for the numbers of (i) evaluation points, that is, of the oracle queries, and (ii) bit-operations involved in our algorithms. Both of the estimates exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
4 Acceleration of subdivision root-finding by means of compression: outline
In Part IV we accelerate the algorithms of both Parts II and III by these factors (of $b$ for roots and $\tilde{b}$ for eigenvalues) to prove the Main Theorems in two ways. Next we outline that acceleration.
(i) Component tree and compact components. Apply subdivision process for a fixed $\sigma$, $1<\sigma<\sqrt{2}$, to Problem 1 and partition the union of all suspect squares tested at every subdivision step into connected components.
Represent all partitions of components with a component tree, $\mathbb{T}$, as follows. At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$. The other (at most $m-1$) vertices of the tree are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between their partitions.
Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(\log(b))$ subdivision iterations, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
Call a component compact if it consists of at most four squares sharing a vertex (see Fig. 2). Recall that a subdivision step never includes new roots into a component and deduce that unless a component stays compact it is soon broken into more than one component or shrinks enough to be covered by an $\epsilon$-disc.
@@ 4 Acceleration of subdivision root-finding by means of compression: outline
(ii) Acceleration of subdivision process with compression of components. Subdivision process becomes quite inefficient where a component stays compact and not broken into more than one component in many consecutive subdivision steps (see Fig. 3). To accelerate it, we replace these steps with compression of such a component into an $\epsilon$-disc or a near-minimal disc $D$ that still keeps all roots lying in the component. In Sec. 19 we comment on two customary but inferior alternatives to our compression, relying on deflation or solely on functional iterations.
@@
If the compression outputs $\epsilon$-disc, we just output it and do not subdivide it anymore. Otherwise we apply subdivision iterations to the minimal superscribing square $S$ of the disc $D$. Since the disc is near-minimal, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc. From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
(iii) Compression of components. We reduce compression of a compact component $C$ to the following two main stages (see more details in Sec. 23.5):
(a) compute a point $c^{\prime}$ lying in or near the minimal disc that covers all roots lying in the component $C$ and let $c^{\prime}$ denote the center of the compressed disc that we seek;
(b) compute a close upper bound $\rho^{\prime}$ on the maximal distance from $c^{\prime}$ to the roots in the component $C$ and output a compressed disc $D(c^{\prime},\rho^{\prime})$.
We prove that Schröder’s single iteration (23.1) following $O(\log(d))$ auxiliary subdivision steps performs step (a) above, that is, computes the desired center $c^{\prime}$.
At the beginning of step (b) above we have a range $[\rho_{-},\rho_{+}]$ for the bound $\rho^{\prime}$, $0<\epsilon\leq\rho_{-}<\rho_{+}\leq 1$. Hence we can approximate it in $\lceil b\rceil$ bisection steps, for $b=\log_{2}(1/\epsilon)$ but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the logarithms, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
We extend our fast soft e/i tests of Part II to fast soft $m$-tests (see part (A) of Sec. 2.1) and use them to define the pivot step of such a modified bisection.
Summarizing we reduce the solution of Problems 0 and 1 to $\tilde{O}(m)$ soft e/i tests and $m$ tests and $\tilde{O}(m)$ Schröder’s iterations. In this way we accelerate the root-finders of both Parts II and III by a factor of $b/\log(b)$ to complete the proof of our Main Theorems.
@@
Many “natural" directions to proving our Main Theorems have led us into dead end. E.g., we defined a tree whose vertices represented non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares. See more examples of this kind in the next remark and Remark 14.1.
@@
Subdivision iterations of the known root-finders, in particular, of those of [24], scale the variable $x$ order of $m$ times by factors of $2^{b}$. Hence the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, which exceeds our bound and that of [96, 103] by a factor of $d$.

### v37

@@ 3.2 Background: error detection and correction
A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc $D$ only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^9] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can compute fast by applying Thm. 12.1 or the algorithm in Sec. 17.5. while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large. We can compute the coefficients of the factor $g(x):=\frac{p(x)}{f(x)}$ for $f(x):=\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $g(x)$ a black box root-finder, and repeat this computation recursively to approximate the $w$ wild roots. [^10] By this recipe we detect any output error at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them. Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them. We readily bound the complexity and error probability of this recursive process (although we will not use these estimates in the following):
@@
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation (cf. example (iii) in Sec. 1.4.1), that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\frac{t^{\prime}(x)}{t(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=w+1}^{m}\frac{1}{x-z_{j}}.$ (3.3)
Remark 3.1.
One can accelerate low precision numerical approximation of the sum $\sum_{j=w+1}^{m}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) by Greengard and Rokhlin [37, 20, 24] (see more in Sec. 24.3).
3.3 Background: root-lifting
Our e/i tests of Part II involve $h$-lifting of roots:
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi\sqrt{-1}}{h}\Big{)},$ (3.4)
@@ 3.3 Background: root-lifting
For $h$ being a power of 2, the $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [38]):
@@
Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.5) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and, due to the map (1.4), from any complex point (see [69, 94, 96] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [96], but even the straightforward implementation of iterations (3.5) promises to be highly efficient as well. The paper [36] explores direct root-squaring for NIR of a black box polynomial:
${\rm NIR}_{p_{h+1}}(x)=\frac{1}{2\sqrt{x}}\left({\rm NIR}_{p_{h}}(\sqrt{x}){\rm NIR}_{p_{h}}(-\sqrt{x})\right),~{}x\neq 0,~{}~{}h=0,1,\dots.$
3.4 Soft e/i test with root-lifting: generic algorithm
In Sec. 7 we elaborate upon the following generic soft e/i test applied to the unit disc $D(0,1)$.
@@ 3.4 Soft e/i test with root-lifting: generic algorithm
INITIALIZATION: Fix two positive integers $h$ and $q$ and $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
COMPUTATIONS: By extending (3.2) to $p_{h}(x)$ of (3.4) (see (6.3)) approximate the values
@@
If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$ ($\sigma$-soft inclusion). Otherwise, claim that #$(D(0,1))=0$ (exclusion).
Remark 3.3.
(6.3) and (3.4) together express $v_{g}$ for all $g$ through $2hq$ values of $p(x)$, by-palink to the values of $p_{h}(s)$ helps us o prove correctness of the algorithm.
@@ 3.5 Soft e/i test with root-lifting: correctness of inclusion claim
It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D\Big{(}c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
@@
We use Cor. 3.1 later but next apply Cor. 3.2 to $p_{h}$ for a sufficiently large $h$ of order $\log(d)$ to prove correctness of the inclusion claim of Alg. 3.1.
@@ 3.6 Soft e/i test with root-lifting: further analysis (outline)
We prove (i) correctness of the exclusion claim of Alg. 3.1 separately for two specifications, both strengthening isolation of the unit circle $C(0,1)$ from the roots by means of root-lifting (3.4).
(a) In Sec. 7.3 we choose any integer $q$ exceeding $d$ and then prove correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Observation 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this e/i test by allowing any $q>m$ (rather than $q>d$) provided that the unit disc contains $m$ roots and $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
@@
(ii) Counting query locations. Eqn. (3.4) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all $h$th roots of $c$. We can fix an integer $h>1$ and apply (3.4) or fix $h>1$ being a power of two and apply (3.4). In both cases Alg. 3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$ when it claims inclusion or at $2qh$ points when it claims exclusion. We always choose $h=O(\log(d))$ and then ignore a factor of $h$ in our $\tilde{O}$ estimates.
In our soft e/i test based on Alg. 3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points and then combine these estimates with Cor. 2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of evaluation queries in the resulting root-finders.
Under the Random Root Model of Sec. 1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$. We strengthen this isolation by applying $O(\log(d))$ root-lifting steps of (3.4) and then whp bound the overall number of queries by $\tilde{O}(bm)$.
(iii) Computational precision and bit-operation complexity. We deduce the same bound $O(b\log(d))$ of [59] on computational precision of our root-finders; our derivation is much simpler because unlike [59] we only use the first order derivatives and divided differences.
We combine this estimate and our count of evaluation queries in our root-finders with (i) Storjohann’s algorithm for Las Vegas computation of the determinant [109] to obtain record fast solution of Problems 2 and $2^{*}$ of matrix eigenvalue approximation (see alternative extension to eigenproblem in part (C) of Sec. 5.4) and (ii) the complexity estimates of Guillaume Moroz [64] for his fast multipoint evaluation of a polynomial represented in monomial basis to arrive at the near-optimal complexity bounds of our Main Theorems 1 – 4 within a factor of $b$, to be removed in Part IV.
3.7 Cauchy-based soft e/i tests: outline
Next we outline our alternative soft e/i test of Part III. Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
@@ 3.7 Cauchy-based soft e/i tests: outline
Having an approximation of integral (3.7) within less than 1/2, we can round it to a closest integer and obtain $s_{0}=\#(D(0,1))$. This root-counter defines $\ell$-tests for all $\ell\leq d$ and in particular an e/i test, for $\ell=1$.
In Part III we approximate the integral (3.7) with finite sums of the values of $\frac{x}{q}$$(x)$ at the scaled $q$th roots of unity. Namely, let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity (cf. (3.4)) and devise randomized $\sigma$-soft e/i tests as follows.
Algorithm 3.2.
[Cauchy e/i test].
(i) Fix a constant $\gamma>1$, say, $\gamma=2$, and choose a real $\rho$ at random in a fixed range in $(1,\sigma)$.
(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
[…]
(iv) Unless $\bar{s}_{0}=0$ certify inclusion.
Otherwise, repeat the test $v$ times for a fixed integer $v>0$ (at the cost of the evaluation of $p(x)$ at $O(mv\log(d))$ points). Unless inclusion is certified in these tests, claim exclusion (correctly with a probability at least $1-1/\gamma^{v}$).
4 Faster subdivision with compression: outline
In Part IV we accelerate the algorithms of both Parts II and III by a factor of $b/\log(b)$ to ensure pairwise isolation of clusters of roots and eigenvalues and to prove Main Theorems 1 – 4; we prove the first three of them in two ways – by devising two distinct supporting root-finders based on Parts II and III, respectively. Next we outline that acceleration.
(i) Component tree and compact components. For fixed integer $m$, positive $\epsilon$, and $\sigma$, $1<\sigma<\sqrt{2}$, apply subdivision process to Problem 1 and at every subdivision step partition the union of the suspect squares into connected components.
@@ 4 Faster subdivision with compression: outline
Represent all partitions of components with a component tree whose at most $m$ leaves represent components covered by isolated $\epsilon$-discs. The other (at most $m-1$) vertices of the tree represent components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between partitions.
Every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(b)$ subdivision iterations, for $b:=\log_{2}(2\Delta^{\prime}/\epsilon)$ and the diameter $\Delta^{\prime}$ of the initial suspect square, every component has diameter below $\epsilon/2$ and hence lies in an $\epsilon$-disc.
(ii) Faster subdivision with compression. Call a component $\mathcal{C}$ compact if all (at most four) its squares share a vertex; subdivision steps fast decrease the diameter of $\mathcal{C}$ but never decrease its distance from external roots (see Fig. 2).
Subdivision becomes inefficient where a component stays unbroken and compact in many subdivision steps (see Fig. 3), and we replace them with its compression with no loss of roots, of which the component very soon becomes strongly isolated.
The compression outputs either an $\epsilon$-disc, and then we stop, or a rigid disc $D$, that is, minimal up to a fixed constant factor, and then we subdivide a minimal superscribing square $S$ of $D$. Since the disc $D$ is rigid, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc.
@@
This way we solve Problem 1 by applying at most $22m-$ 2 compression steps and in addition $O(m)$ soft e/i tests, as we deduced in Sec. 19.
(iii) Compression of components. We compress a compact component $\mathcal{C}$ into a disc $D=D(c,\rho)$ by computing:
(a) the number of roots in a component,
(b) a point $c$ in or near the minimal disc covering the root set of $\mathcal{C}$ (hereafter the MCD of $\mathcal{C}$) and
(c) a reasonably close upper bound $\rho$ on the maximal distance from $c$ to the roots in $\mathcal{C}$.
In Secs. 17.4 and 17.5 we readily perform steps (a) and (b) at a low computational cost because subdivision iterations fast strengthen the isolation of the component from external roots.
We begin step (c) with a range $[\rho_{-},\rho_{+}]$ for the bound $\rho$, $0<1/2^{b}=\rho_{-}<\rho_{+}$. We can approximate $\rho$ by applying bisection to this range but proceed faster by applying bisection of the logarithms, said to be BoL, to the range $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
We apply fast soft $m$-tests (see part (A) of Sec. 2.1) to define a pivot step of our accelerated bisection, with which we ultimately reduce compression to $O(m\log(b))$ e/i tests. Their computational cost dominates the overall cost of the solution of Problems 0,0, 1, and 1, and this enables us to complete the proof of our Main Theorems.
@@
Many “natural" directions to proving our Main Theorems have led us into dead end. E.g., we tried to use a tree whose vertices represented components made up of non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares. See more examples of this kind in the next remark and Remark 7.1.
@@
Subdivision iterations of the known root-finders, in particular, of those of [17], scale the variable $x$ order of $m$ times by factors of $2^{b}$. Hence the bit-complexity specified in the Main Theorem of [17] has order of $mbd^{2}$, which exceeds our bound and that of [78, 84] by a factor of $d$ (see more comments in Sec. 10.4).